# How hydrodynamic interactions alter polymer stretching in turbulence

### v1

@@ IV.1 HI steepen the coil-stretch transition
We expect HI to alter the relaxation time of a chain, in a manner that depends on the number of beads. This implies that the true Weissenberg number of a chain with HI should differ from $\mathrm{Wi}=\lambda\tau^{D}$, the value for a free draining chain, by a factor that depends on $h^{\star}$ and $N_{\mathrm{b}}$. Now, if such a rescaling of the Weissenberg number was the only effect of HI, then the results in Fig. 1 should show that increasing $h^{\star}$ simply translates the curves of $\langle R\rangle$ versus $\mathrm{Wi}$ to the left (for dumbbells) or to the right (for chains). We now check if this is the case. We first define a critical $\mathrm{Wi}_{\rm c}$ as the value of $\mathrm{Wi}$ at which $\langle R\rangle$ attains half its large-$\mathrm{Wi}$ asymptote. We then define a rescaled $\mathrm{Wi}^{\star}\equiv(1/2)\mathrm{Wi}/\mathrm{Wi}_{\rm c}$, so that the curves for all $h^{\star}$ coincide at the transition point $\mathrm{Wi}^{\star}=1/2$. The question is whether the curves entirely collapse after this rescaling or not.
Figure 3: Mean extension $\langle R\rangle/R_{eq}$ for various values of the rescaled Weissenberg number $\mathrm{Wi}^{\star}$, considering $h^{\star}$ = 0, 0.2, 0.49 [see the legend in panel (b)]: Results are presented for (a) dumbbells, (b) twenty-bead chains. (c) four-bead chains, and (d) ten-bead chains. The inset of panel (a) is a zoom of the main panel. The three curves of $\langle R\rangle/R_{eq}$ in panel (c) almost entirely overlap. The inset of panel (c) shows the critical value $\mathrm{Wi}_{c}$ of the coil-stretch transition used to rescale $\mathrm{Wi}$, so that $\mathrm{Wi}^{\star}_{\mathrm{c}}=1/2$ for all $h^{\star}$.
@@
The rescaled plots of $\langle R\rangle$ as a function of $\mathrm{Wi}^{\star}$ are presented in Fig. 3. The rescaling accounts quite well for the weak effect of HI on dumbbells [Fig. 3(a)]. The effect on chains [Fig. 3(b)], however, is subtle and cannot be accounted for by rescaling $\mathrm{Wi}$, i.e, by using a HI-modified elastic relaxation time. Instead, we see that HI steepens the coil-stretch transition [Fig. 3(b)], causing chains below $\mathrm{Wi}^{\star}_{\mathrm{c}}$ to stretch less and those above $\mathrm{Wi}^{\star}_{\mathrm{c}}$ to stretch more. This effect is stronger for chains with more beads [compare the results for $N_{\mathrm{b}}=20$ and 10 in Fig. 3(b,d)]. Indeed, for a given non-zero value of $h^{\star}$, chains with more beads undergo a sharper coil-stretch transition (see Fig. S1 of the supplemental material [69], where the results of Fig. 3 are replotted to facilitate a comparison of chains with the same $h^{\star}$ but different $N_{\mathrm{b}}$).
@@
We have constructed the rescaling to give $\mathrm{Wi}^{\star}=1/2$ because the value 1/2 marks the transition in the theory of Balkovsky et al. [34, 35]. In their work, however, the transition was described in terms of the exponent of the power-law scaling of the PDF of $R$. We cannot follow the same approach here because, as we show in the next section, the PDF is qualitatively modified by HI to the extent that a clear power-law range is no longer visible. We have therefore used a criterion based on $\langle R\rangle/R_{\mathrm{eq}}$, which yields a $\mathrm{Wi}_{\rm c}$ that differs from 1/2 even for free-draining chains [see the inset of Fig. 3(c)]. More important than the precise values of $\mathrm{Wi}_{\rm c}$, which depend on the transition criterion, is the variation of $\mathrm{Wi}_{\rm c}$ with $h^{\star}$ shown in the inset of Fig. 3(c). $\mathrm{Wi}_{\rm c}$ decreases with $h^{\star}$ for dumbbells and increases with $h^{\star}$ for multibead chains, which implies that the effective relaxation time of a chain (with a fixed spring time-scale $\tau_{s}$) increases with $h^{\star}$ for dumbbells and decreases with $h^{\star}$ for chains. This dependence of the effective relaxation time on HI is captured by the rescaled Weissenberg number, $\mathrm{Wi}^{\star}$. So, henceforth, we present results in terms of $\mathrm{Wi}^{\star}$.
@@ IV.2 HI limit self-similar stretching
The PDF of $R$ for dumbbells is presented in Fig. 4(a), for $\mathrm{Wi}^{\star}$ = 0.3, 0.5, and 2.0, both with and without HI (solid lines are for $h^{\star}=0.2$ while the dashed lines are for $h^{\star}=0$). The logarithmic axes of this plot makes apparent the power-law behaviour of the PDF for $R_{\mathrm{eq}}\lesssim R\lesssim R_{\mathrm{m}}$. Importantly, we see that HI does not disrupt the power-law behaviour, though it does increase slightly the power-law exponent, so that the fraction of shrunk dumbbells ($R\sim R_{\mathrm{eq}}$) is marginally reduced in favour of stretched ones ($R\gg R_{\mathrm{eq}}$). This result is consistent with recent analytical predictions for HI-endowed dumbbells in a random chaotic flow [24].
@@
The PDF of $R$ for chains is affected quite differently by HI, as shown by Fig. 4(b) which presents results for twenty-bead chains. We see a substantial increase in the fraction of coiled polymers ($R\sim R_{\mathrm{eq}}$) and a corresponding reduction in the stretched polymers ($R\gg R_{\mathrm{eq}}$). This HI-induced modification changes the shape of the PDF, for small and intermediate values of $\mathrm{Wi}^{\star}$, and limits the power-law range. Taking the case of $\mathrm{Wi}^{\star}=0.3$ as an example [red lines in Fig. 4(b)], we see that the power-law variation, which in the absence of HI held sway for $R_{\mathrm{eq}}\lesssim R\lesssim R_{\mathrm{m}}$, emerges in the presence of HI only for $20R_{\mathrm{eq}}\lesssim R\lesssim R_{\mathrm{m}}$. For $R>20R_{\mathrm{eq}}$, the beads of the $N_{\mathrm{b}}=20$ chain are sufficiently far apart for the strength of HI to diminish. For large $\mathrm{Wi}^{\star}$ almost all chains (and dumbbells) are strongly stretched and nearly unaffected by HI [see the PDF for $\mathrm{Wi}^{\star}=2.0$ in Figs. 4(a,b)].
@@
The effects of HI increase in magnitude with $h^{\star}$ while remaining qualitatively the same [see Fig. S2(c) in the supplemental material [69] where the analogue of Fig. 4(b) is presented for $h^{\star}=0.49$]. Also, as already demonstrated by the results for $\langle R\rangle$ in Fig. 3(b,d), reducing the number of beads in the chain weakens the hydrodynamic shielding effect of HI. So, the changes to the PDF of $R$ are similar but smaller for ten-bead chains, when compared to twenty-bead chains [see Fig. S2(d) in the supplemental material [69]].
@@
Figure 5 shows how the PDF of $R$ changes on increasing the number of beads, while keeping $h^{\star}$ fixed. We see that the larger the value of $h^{\star}$, the greater are the changes produced by including more beads in the chain [compare Figs. 5(a,b) with Figs. 5(c,d)]. It is natural to ask how much further the stretching behavior of chains would change if one were to continue increasing $N_{\mathrm{b}}$ up to the number of Kuhn steps $N_{\mathrm{k}}$. This question—about the asymptotic dependence on the number of beads of the dynamics of a chain with HI—is key to the development of successive fine graining schemes for accurately modeling polymers [70]; it has been well-studied in the context of equilibrium properties [7] and viscometric flows [67, 71, 70, 72, 73]. To answer this question in turbulent flows would require increasing $N_{\mathrm{b}}$ well beyond twenty, a computationally demanding and important task for future work.
@@ IV.3 HI delay migrations between coiled and stretched states
Now in turbulence, the strong fluctuations of the velocity gradient cause polymers to sample a wide range of extensions and to migrate between states of large and small extension relatively quickly. The evolution of the end-to-end extension is illustrated in Fig. 7, which presents a typical time-trace of $R(t)$ for a twenty-bead chain ($\mathrm{Wi}^{\star}=0.5$), with and without HI (the same Lagrangian trajectory, and hence the same time series for $\bm{\kappa}(t)$, is used for both cases). We see that, even with HI, chains repeatedly stretch and recoil as they move through the turbulent flow. Therefore, after an initial transient (which lasts less than $\sim 40\lambda t$ for the range of $\mathrm{Wi}^{\star}$ considered here), a stationary regime is attained whose statistics are independent of the initial configuration. Hence, HI cannot induce a coil-stretch hysteresis in the extension statistics measured in turbulence [38]. However, it is possible that HI modify the typical time it takes for chains to migrate between states of small and large extension. Indeed, Fig. 7 shows that the chain with HI attains both smaller and larger extensions than the chain without HI. The HI-endowed chain experiences a prolonged episode with very small extensions. However, when both chains are stretched by the flow, the fluctuations towards smaller extensions seem to be weaker for the HI-endowed chain. To quantitatively investigate the migration between small and large extensions, we define two thresholds, $\ell_{\mathrm{coil}}$ and $\ell_{\mathrm{str}}$, and designate chains with $R<\ell_{\mathrm{coil}}$ as coiled and those with $R>\ell_{\mathrm{str}}$ as stretched. We then compute the intervals of time for which chains remain in these two states, and thereby obtain the PDF of persistence times in both states. We use $\ell_{\mathrm{coil}}=5R_{\mathrm{eq}}$ and $\ell_{\mathrm{str}}=50R_{\mathrm{eq}}$, after checking that our conclusions do not depend on the precise values of these thresholds. A similar analysis of the persistence time in stretched states was performed for Rouse chains (no HI) in Picardo et al. [39], where the corresponding PDF was found to have an exponential tail.
@@
Figures 7(a-c) present the PDFs of the persistence time in coiled states, $t_{\mathrm{coil}}$, for twenty-bead chains with (a) $\mathrm{Wi}^{\star}=0.3$, (b) $\mathrm{Wi}^{\star}=0.5$, and (c) $\mathrm{Wi}^{\star}=1.0$. The exponential-tailed distributions are seen to shift significantly towards larger values of $t_{\mathrm{coil}}$ as $h^{\star}$ is increased. The same is true for the PDFs of the persistence time in stretched states, $t_{\mathrm{str}}$, as seen in Figs. 7(d-f). Clearly, HI does increase the time for migrating between coiled and stretched states. This effect is inline with HI inducing an effective conformation-dependent drag.
@@
In Fig. S3 of the supplemental material, we present the persistence time PDFs for ten-bead chains and for dumbbells [69]. As expected from the discussion in Sec. III, the results for dumbbells are opposite to those for chains, with HI-endowed dumbbells taking less time to migrate between shrunk and stretched states. The results for ten-bead chains are similar to those for twenty-bead chains, but with weaker effects of HI owing to the reduction of the number of beads.

### v2

@@ IV.1 HI steepen the coil-stretch transition
We expect HI to alter the relaxation time of a chain, in a manner that depends on the number of beads. This implies that the true Weissenberg number of a chain with HI should differ from $\mathrm{Wi}=\lambda\tau^{D}$, the value for a free draining chain, by a factor that depends on $h^{\star}$ and $N_{\mathrm{b}}$. Now, if such a rescaling of the Weissenberg number was the only effect of HI, then the results in Fig. 1 should show that increasing $h^{\star}$ simply translates the curves of $\langle R\rangle$ versus $\mathrm{Wi}$ to the left (for dumbbells) or to the right (for chains). We now check if this is the case. We first define a critical $\mathrm{Wi}_{\rm c}$ as the value of $\mathrm{Wi}$ at which $\langle R\rangle$ attains half its large-$\mathrm{Wi}$ asymptote. We then define a rescaled $\mathrm{Wi}^{\star}\equiv(1/2)\mathrm{Wi}/\mathrm{Wi}_{\rm c}$, so that the curves for all $h^{\star}$ coincide at the transition point $\mathrm{Wi}^{\star}=\mathrm{Wi}^{\star}_{c}=1/2$. This rescaling is therefore an empirical shift of the Weissenberg number, based on a $h^{\star}$ dependent time scale that is determined from the results, post facto, such that the mean extension undergoes a rapid increase at $\mathrm{Wi}^{\star}=1/2$ for all values of $h^{\star}$. One may view $\mathrm{Wi}^{\star}$ as the Weissenberg number that would be obtained if the $h^{\star}$ dependent time scale that controls the coil-stretch transition were used to nondimensionalize the Lyapunov exponent. The question is whether the curves entirely collapse after this rescaling or not.
@@
The rescaled plots of $\langle R\rangle$ as a function of $\mathrm{Wi}^{\star}$ are presented in Fig. 3. For dumbbells, the weak effects of HI, seen in Fig. 1(a), become negligible after the rescaling [see Fig. 3(a)]. In contrast, the much stronger effects of HI on twenty-bead chains [Fig. 1(b)] cannot be accounted for by rescaling $\mathrm{Wi}$, i.e, by using a HI-modified elastic relaxation time. Indeed, Fig. 3(b) shows that HI steepens the coil-stretch transition, by causing chains below $\mathrm{Wi}^{\star}_{\mathrm{c}}$ to stretch less and those above $\mathrm{Wi}^{\star}_{\mathrm{c}}$ to stretch more. In other words, on comparing chains with and without HI, at the same distance from the corresponding coil-stretch transitions, we find that HI causes stiff chains to stretch less and moderately elastic chains to stretch more.
@@
The inset of Fig. 3(c), which presents the values of $\mathrm{Wi}_{c}$ for different $N_{\mathrm{b}}$ and $h^{\star}$, shows how HI shifts the coil-stretch transition for chains with different numbers of beads. Importantly, the shift changes direction and increases in magnitude as $N_{\mathrm{b}}$ is increased from 2 to 20. Thus, while $\mathrm{Wi}_{\rm c}$ decreases slightly with $h^{\star}$ for dumbbells (compare $\mathrm{Wi}_{c}$ for $h^{\star}=0$ and 0.49, at $N_{\mathrm{b}}=2$), it increases strongly with $h^{\star}$ for twenty bead chains ($N_{\mathrm{b}}=20$). (Note that we have used $\mathrm{Wi}_{c}$ to rescale $\mathrm{Wi}$ such that $\mathrm{Wi}^{\star}_{c}=1/2$ at the transition because the value 1/2 marks the transition in the theory of Balkovsky et al. [37, 38]. In their work, however, the transition was described in terms of the exponent of the power-law scaling of the PDF of $R$. We cannot follow the same approach here because, as we show in the next section, the PDF is qualitatively modified by HI to the extent that a clear power-law range is no longer visible. We have therefore used a criterion based on $\langle R\rangle/R_{\mathrm{eq}}$, which yields a $\mathrm{Wi}_{\rm c}$ that differs from 1/2 even for free-draining chains [see the inset of Fig. 3(c)].)
The results in this section show that one cannot capture the effects of HI on a multibead chain by adding HI to a dumbbell. Being unable to form a physical coil, the dumbbell cannot represent the effects of hydrodynamic shielding, and so its behavior in the presence of HI departs from that of chains. The disagreement is both quantitative and qualitative: not only is the effect of HI on the extension of dumbbells much weaker than that on chains, but it also acts in the opposite direction. This contrast is evident in the variation of the location of the coil-stretch transition $\mathrm{Wi}_{c}$, which shifts slightly towards smaller $\mathrm{Wi}$ for dumbbells but strongly towards larger $\mathrm{Wi}$ for chains (inset of Fig. 3(c)). Furthermore, even after accounting for this difference in the shift of the transition, by rescaling $\mathrm{Wi}$ to $\mathrm{Wi}^{\star}$, chains with HI exhibit a steepening of the transition that is absent for dumbbells [compare Fig. 3(a) and Fig. 3(b)].
Hence, one has to refine the dumbbell model by adding more beads in order to obtain, even qualitatively, the behavior of chains with HI. Quantitatively, we find, for the limited range of $N_{\mathrm{b}}$ considered here, that the effects of HI increase systematically as the number of beads are increased (i.e., as the model is refined for fixed values of the polymer’s equilibrium and contour lengths, $R_{\mathrm{m}}$ and $R_{\mathrm{eq}}$; see Sec. II.2). Fig. 3(d) shows that the effects of HI are much weaker in ten-bead chains when compared to twenty-bead chains [Fig. 3(b)]. In fact, for a given non-zero value of $h^{\star}$, chains with more beads undergo a sharper coil-stretch transition (see Fig. S1 of the supplemental material (SM) [76], where the results of Fig. 3 are replotted to facilitate a comparison of chains with the same $h^{\star}$ but different $N_{\mathrm{b}}$).
Figure 3: Mean rescaled extension $\langle R\rangle/R_{eq}$ for various values of the rescaled Weissenberg number $\mathrm{Wi}^{\star}$, considering $h^{\star}$ = 0, 0.2, 0.49 [see the legend in panel (b)]: Results are presented for (a) dumbbells, (b) twenty-bead chains. (c) four-bead chains, and (d) ten-bead chains. The inset of panel (a) is a zoom of the main panel. The three curves of $\langle R\rangle/R_{eq}$ in panel (c) almost entirely overlap. The inset of panel (c) shows the critical value $\mathrm{Wi}_{c}$ of the coil-stretch transition used to rescale $\mathrm{Wi}$, so that $\mathrm{Wi}^{\star}_{\mathrm{c}}=1/2$ for all $h^{\star}$.
Having examined how the coil-stretch transition is affected by HI, we proceed in the following sections to compare the statistics of extension of chains with different $h^{\star}$ and $N_{\mathrm{b}}$ at the same distance from their corresponding transition points. To facilitate this comparison, which reveals the effects of HI that cannot be accounted for by a $h^{\star}$-dependent time-scale, we henceforth present results in terms of $\mathrm{Wi}^{\star}$.
@@ IV.2 HI limit self-similar stretching
The PDF of $R$ for dumbbells is presented in Fig. 4(a), for $\mathrm{Wi}^{\star}$ = 0.3, 0.5, and 2.0, both with and without HI (solid lines are for $h^{\star}=0.2$ while the dashed lines are for $h^{\star}=0$). The logarithmic axes of this plot makes apparent the power-law behaviour of the PDF for $R_{\mathrm{eq}}\lesssim R\lesssim R_{\mathrm{m}}$. Importantly, we see that HI does not disrupt the power-law behaviour, though, for small and intermediate $\mathrm{Wi}^{\star}$, it does increase slightly the power-law exponent, so that the fraction of shrunk dumbbells ($R\sim R_{\mathrm{eq}}$) is marginally reduced in favour of stretched ones ($R\gg R_{\mathrm{eq}}$). This result is consistent with recent analytical predictions for HI-endowed dumbbells in a random chaotic flow [17].
@@
The results for chains are again very different from that for dumbbells. Fig. 4(b), which presents the PDF of $R$ for twenty-bead chains, shows that HI produce a substantial increase in the fraction of coiled polymers ($R\sim R_{\mathrm{eq}}$) and a corresponding reduction in the fraction of stretched polymers ($R\gg R_{\mathrm{eq}}$). This HI-induced modification changes the shape of the PDF, for small and intermediate values of $\mathrm{Wi}^{\star}$, and limits the power-law range. Taking the case of $\mathrm{Wi}^{\star}=0.3$ as an example [red lines in Fig. 4(b)], we see that the power-law variation, which in the absence of HI held sway for $R_{\mathrm{eq}}\lesssim R\lesssim R_{\mathrm{m}}$, emerges in the presence of HI only for $20R_{\mathrm{eq}}\lesssim R\lesssim R_{\mathrm{m}}$. For $R>20R_{\mathrm{eq}}$, the beads of the $N_{\mathrm{b}}=20$ chain are sufficiently far apart for the strength of HI to diminish. For large $\mathrm{Wi}^{\star}$ almost all chains (and dumbbells) are strongly stretched and nearly unaffected by HI [see the PDF for $\mathrm{Wi}^{\star}=2.0$ in Figs. 4(a,b)].
Figure 5: (a,c) Variance $\Sigma=\langle(R-\langle R\rangle)^{2}\rangle/R_{\mathrm{eq}}^{2}$ and (b,d) skewness $\Gamma=\langle(R-\langle R\rangle)^{3}\rangle/\Sigma^{(3/2)}$ of the PDF of $R$, for various values of the rescaled Weissenberg number $\mathrm{Wi}^{\star}$, considering $h^{\star}$ = 0, 0.2, 0.49 [see the legend in panel (a)]. Results for twenty bead chains and dumbbells are presented in the top and bottom rows, respectively. The corresponding PDFs for $\mathrm{Wi}^{\star}=0.3,$ 0.5, and 2.0 are shown in Fig. 4(b) and Fig. 4(a).
@@
The effects of HI increase in magnitude with $h^{\star}$ while remaining qualitatively the same [see Fig. S2(c) in the SM [76] where the analogue of Fig. 4(b) is presented for $h^{\star}=0.49$]. Also, as already demonstrated by the results for $\langle R\rangle$ in Fig. 3(b,d), reducing the number of beads in the chain weakens the hydrodynamic shielding effect of HI. So, the changes to the PDF of $R$ are similar but smaller for ten-bead chains, when compared to twenty-bead chains [see Fig. S2(d) in the SM [76]].
Figures 5(a,b) quantify the influence of HI on the PDF of $R$, for twenty bead chains, by plotting (a) the variance $\Sigma=\langle(R-\langle R\rangle)^{2}\rangle/R_{\mathrm{eq}}^{2}$ and (b) the skewness $\Gamma=\langle(R-\langle R\rangle)^{3}\rangle/\Sigma^{(3/2)}$ as function of $\mathrm{Wi}^{\star}$ for different values of $h^{\star}$. In the absence of HI, the variance $\Sigma$ is known to peak near the coil-stretch transition, $\mathrm{Wi}^{\star}=1/2$, where the distribution is at its broadest [40, 45, 48]. We see the same behavior of $\Sigma$ in the presence of HI; the peak at $\mathrm{Wi}^{\star}=1/2$ becomes more prominent as $h^{\star}$ increases because the corresponding PDF develops a bimodal shape, with local peaks near $R_{\mathrm{eq}}$ and $R_{\mathrm{m}}$ [see the dashed and solid green lines in Fig. 4(b), as well as the blue lines in Figs. 6(b) and 6(d) below]. The restriction by HI of the power-law range of the PDF of $R$ for stiff chains, noted earlier, is associated with an increase in the skewness of the distribution [see the red curves for $\mathrm{Wi}^{\star}=0.3$ in Fig. 4(b)]; Fig. 5(b) quantifies the extent of this effect showing a strong increase in $\Gamma$ with $h^{\star}$ for small values of $\mathrm{Wi}^{\star}$.
In the case of dumbbells [Figs. 5(c,d)], $\Sigma$ and $\Gamma$ exhibit a much weaker dependence on HI as compared to twenty bead chains; moreover, the changes occur in the opposite direction, so that $\Sigma$ and $\Gamma$ decrease slightly with $h^{\star}$ in the case of dumbbells [see the insets of Figs. 5(c,d)], whereas they increase strongly with $h^{\star}$ in the case of twenty-bead chains [Figs. 5(a,b)]. Thus, adding HI to a dumbbell fails to capture the effects of HI on the extension of multibead chains, with regard to not only the mean [Fig. 3(a,b)] but also the distribution.
@@
Figure 6 depicts how the PDF of $R$ changes on increasing the number of beads, while keeping $h^{\star}$ fixed. The PDF changes systematically with $N_{\mathrm{b}}$, such that the power-law distribution for dumbbells ($N_{\mathrm{b}}=2$) is altered substantially once $N_{\mathrm{b}}$ reaches 20. At small $\mathrm{Wi}^{\star}$, increasing $N_{\mathrm{b}}$ produces an increase in the fraction of small-extension configurations [Figs. 6(a,c)] due to the action of hydrodynamic shielding which grows in relevance with $N_{\mathrm{b}}$. As a consequence, the power-law behavior is progressively disrupted and the skewness of the PDF increases; indeed, $\Gamma$ in Figs. 5(b,d) is much larger for $N_{\mathrm{b}}=20$ than for $N_{\mathrm{b}}=2$, at small $\mathrm{Wi}^{\star}$. Near the coil-stretch transition, the PDF develops a bimodal character as $N_{\mathrm{b}}$ is increased [Figs. 6(b,d)], which in turn increases the variance; thus, $\Sigma$ in Figs. 5(a,c) is much larger for $N_{\mathrm{b}}=20$ than for $N_{\mathrm{b}}=2$, for $\mathrm{Wi}^{\star}\approx 1/2$. The extent of these changes with $N_{\mathrm{b}}$ is greater for larger values of $h^{\star}$ [compare Figs. 6(a,b) with Figs. 6(c,d)].
It is natural to ask how much further the stretching behavior of chains would change if one were to continue increasing $N_{\mathrm{b}}$ up to the number of Kuhn steps $N_{\textsc{k}}$. This question—about the asymptotic dependence on the number of beads of the dynamics of a chain with HI—is key to the development of successive fine graining schemes for accurately modeling polymers [52]; it has been well-studied in the context of equilibrium properties [7] and viscometric flows [65, 77, 52, 78, 79]. To answer this question in turbulent flows would require increasing $N_{\mathrm{b}}$ well beyond twenty, a computationally demanding and important task for future work. In the absence of a fine-graining scheme for chains with HI in turbulence, we do not know how to adjust $h^{\star}$, as $N_{\mathrm{b}}$ is increased, so that the different chains correspond to increasingly refined models of the same polymer. We have therefore kept $h^{\star}$ fixed and varied $N_{\mathrm{b}}$ in Fig. 6 to show how the sensitivity of the extensional dynamics to HI increases with the number of beads.
@@ IV.3 HI delay migrations between coiled and stretched states
Now in turbulence, the strong fluctuations of the velocity gradient cause polymers to sample a wide range of extensions and to migrate between states of large and small extension relatively quickly. The evolution of the end-to-end extension is illustrated in Fig. 8, which presents a typical time-trace of $R(t)$ for a twenty-bead chain ($\mathrm{Wi}^{\star}=0.5$), with and without HI (the same Lagrangian trajectory, and hence the same time series for $\bm{\kappa}(t)$, is used for both cases). We see that, even with HI, chains repeatedly stretch and recoil as they move through the turbulent flow. Therefore, after an initial transient (which lasts less than $\sim 40\lambda t$ for the range of $\mathrm{Wi}^{\star}$ considered here), a stationary regime is attained whose statistics are independent of the initial configuration. Hence, HI cannot induce a coil-stretch hysteresis in the extension statistics measured in turbulence [41]. However, it is possible that HI modify the typical time it takes for chains to migrate between states of small and large extension. Indeed, Fig. 8 shows that the chain with HI attains both smaller and larger extensions than the chain without HI. The HI-endowed chain experiences a prolonged episode with very small extensions. However, when both chains are stretched by the flow, the fluctuations towards smaller extensions seem to be weaker for the HI-endowed chain.
To quantitatively investigate the migration between small and large extensions, we define two thresholds, $\ell_{\mathrm{coil}}$ and $\ell_{\mathrm{str}}$, and designate chains with $R<\ell_{\mathrm{coil}}$ as coiled and those with $R>\ell_{\mathrm{str}}$ as stretched. We then compute the intervals of time for which chains remain in these two states, and thereby obtain the PDF of persistence times in both states. We use $\ell_{\mathrm{coil}}=5R_{\mathrm{eq}}$ and $\ell_{\mathrm{str}}=50R_{\mathrm{eq}}$, after checking that our conclusions do not depend on the precise values of these thresholds [see the SM [76]]. A similar analysis of the persistence time in stretched states was performed for Rouse chains (no HI) in Picardo et al. [42], where the corresponding PDF was found to have an exponential tail.
@@
Figures 8(a-c) present the PDFs of the persistence time in coiled states, $t_{\mathrm{coil}}$, for twenty-bead chains with (a) $\mathrm{Wi}^{\star}=0.3$, (b) $\mathrm{Wi}^{\star}=0.5$, and (c) $\mathrm{Wi}^{\star}=1.0$. In the presence of HI, the PDF continues to have an exponential tail and broadens so that the probability of experiencing larger values of $t_{\mathrm{coil}}$ increases significantly as $h^{\star}$ is increased. The same is true for the PDFs of the persistence time in stretched states, $t_{\mathrm{str}}$, as seen in Figs. 8(d-f). Clearly, HI does increase the time for migrating between coiled and stretched states. This effect is inline with HI inducing an effective conformation-dependent drag.
@@
In Fig. S3 of the SM, we present the persistence time PDFs for ten-bead chains and for dumbbells [76]. As expected from the discussion in Sec. III, the results for dumbbells are opposite to those for chains, with HI-endowed dumbbells taking less time to migrate between shrunk and stretched states. The results for ten-bead chains are similar to those for twenty-bead chains, but with weaker effects of HI owing to the reduction of the number of beads.